# Symmetry-aware spectral bounds for the electronic structure Hamiltonian

### v1

@@ C.2 Application: electronic structure Hamiltonian
Appendix D Numerical evaluation of symmetry-aware spectral bounds
The numerical evaluation of the spectral bounds requires finding the maximum and minimum energy eigenvalues of a particular Hermitian operator ($\hat{H}_{\mathrm{1body}}$, $\hat{H}_{\mathrm{2body}}$ or $\hat{H}$) in every symmetry sector $\mu$. As alluded in the main text, we consider the total particle number $\hat{N}$, total spin $\hat{S}^{2}$, and spin-projection $\hat{S}_{z}$ symmetries which can be uniquely identified by the set of quantum numbers, $\mu=(\eta,s,m_{s})$. Here, $\eta$ is equal to the total number of electrons, $s$ is the spin quantum number, and $m_{s}$ is the spin-projection number, ranging from -$s$ to $s$. In a fixed basis set with $N_{\mathrm{orb}}$ spatial orbitals, the total number of electrons will be limited to the range $[0,2N_{\mathrm{orb}}]$ across the entire Hilbert space. In the absence of external magnetic fields or spin-orbit coupling, the latter resulting from relativistic adjustments to the Hamiltonian, the energy eigenvalues characterized by constant $\eta$ and $s$, but varying $m_{s}$, will be degenerate. This implies that the total number of symmetry sectors $(\eta,s)$ with unique energy eigenvalues will be given by,
$\frac{(N_{\mathrm{orb}}+1)(N_{\mathrm{orb}}+2)}{2}.$ (C.50)
While this number is only quadratic, it renders the numerical evaluation of $\Delta/2$ and $\Delta_{s}/2$ quite resource-intensive for large basis sets. In the following, we discuss a numerically tractable way of evaluating the maximum and minimum eigenvalues based on finding the orbital-optimized determinant that either maximizes or minimizes the energy, akin to a restricted Hartree-Fock (closed shell RHF and restricted open-shell ROHF) calculation.
D.1 Orbital Optimization Procedure
$E_{\mathrm{HF}}(\boldsymbol{\kappa})\!=\!\sum_{i}^{\eta_{\alpha}}h_{ii}(\boldsymbol{\kappa})+\sum_{i}^{\eta_{\beta}}h_{ii}(\boldsymbol{\kappa})+\!\frac{1}{2}\!\!\left(\!\sum_{i,j}^{\eta_{\alpha},\eta_{\alpha}}\!\!g_{iijj}(\boldsymbol{\kappa})+\!\sum_{i,j}^{\eta_{\beta},\eta_{\beta}}\!\!g_{iijj}(\boldsymbol{\kappa})+2\!\!\sum_{i,j}^{\eta_{\alpha},\eta_{\beta}}\!\!g_{iijj}(\boldsymbol{\kappa})-\!\sum_{i,j}^{\eta_{\alpha},\eta_{\alpha}}\!\!g_{ijji}(\boldsymbol{\kappa})-\!\sum_{i,j}^{\eta_{\beta},\eta_{\beta}}\!\!g_{ijji}(\boldsymbol{\kappa})\right)\!\!,$ (C.51)
with orbital-transformed one-electron $h_{pq}$ and two-electron $g_{pqrs}$ tensor coefficients,
$h_{pq}(\boldsymbol{\kappa})$ $=\sum_{ij}h_{ij}[\exp(-\boldsymbol{\kappa})]_{ip}[\exp(-\boldsymbol{\kappa})]_{jq},$ (C.52)
$g_{pqrs}(\boldsymbol{\kappa})$ $=\sum_{ijkl}g_{ijkl}[\exp(-\boldsymbol{\kappa})]_{ip}[\exp(-\boldsymbol{\kappa})]_{jq}[\exp(-\boldsymbol{\kappa})]_{kr}[\exp(-\boldsymbol{\kappa})]_{ls},$ (C.53)
defined with respect to the unitary matrix, $\exp(-\boldsymbol{\kappa})$, written in terms of the anti-Hermitian matrix, $\boldsymbol{\kappa}$, where $\boldsymbol{\kappa}^{\dagger}=-\boldsymbol{\kappa}$. To improve the stability of the optimization procedure (especially relevant for finding the maximun eigenvalue), we used the pre-orthogonalized one-electron $h_{ij}$ and two-electron integrals $g_{ijkl}$,
$h_{ij}$ $=\sum_{\mu\nu}h_{\mu\nu}[S^{-\frac{1}{2}}]_{\mu i}[S^{-\frac{1}{2}}]_{\nu j},$ (C.54)
$g_{ijkl}$ $=\sum_{\mu\nu\rho\sigma}g_{\mu\nu\rho\sigma}[S^{-\frac{1}{2}}]_{\mu i}[S^{-\frac{1}{2}}]_{\nu j}[S^{-\frac{1}{2}}]_{\rho j}[S^{-\frac{1}{2}}]_{\sigma l},$ (C.55)
as our initial starting guess. These quantities are defined with respect the atomic orbital overlap matrix ($S_{\mu p}$) as well as the one-electron ($h_{\mu\nu}$) and two-electron ($g_{\mu\nu\rho\sigma}$) integrals defined in the atomic orbital basis. For the optimization of $N$ real spatial orbitals, it is sufficient to consider the real part of $\boldsymbol{\kappa}$, which contains only $N_{\mathrm{orb}}(N_{\mathrm{orb}}-1)/2$ parameters. The maximum eigenvalue, $E_{\mathrm{max}}^{(\eta_{\alpha},\eta_{\beta})}$, can be found with the same numerical minimizer by taking the negative of the one-electron and two-electron integrals. It is important to note that for RHF and ROHF techniques, the associated wave function will be an eigenfunction of the total spin operator, $\hat{S}^{2}$, with an expectation value equal to,
$\braket{\hat{S}^{2}}=\left(\frac{\eta_{\alpha}-\eta_{\beta}}{2}\right)\!\!\left(\frac{\eta_{\alpha}-\eta_{\beta}}{2}+1\right),$ (C.56)
assuming $\eta_{\alpha}\geq\eta_{\beta}$. As a result, the spin quantum number $s$ is equal to the the spin-projection quantum number $|m_{s}|$. Although this might appear restrictive, this implies that in scenarios devoid of magnetic fields and spin-orbit interactions, adjusting $\eta_{\alpha}$ and $\eta_{\beta}$ suffices to encompass all possible symmetry sectors characterized by unique energy eigenvalues.
D.2 Numerical limitations
@@
While the proposed numerical method provides a scalable approach to compute symmetry-aware spectral bounds for large-scale systems, it does have important limitations and challenges which we discuss below. First, as alluded to above, we found that the numerical procedure was prone to certain convergence issues arising from the highly non-convex landscape encountered in the orbital optimization procedure. In particular, we found that random initial guesses highly affected that final converged energy values. To address this issue, we proposed the pre-orthogonalized one-electron and two-electron integrals in Eqs. (D.5) and (D.6). We also found that initializing the $N_{\mathrm{orb}}(N_{\mathrm{orb}}-1)/2$ parameters to zero helped in obtaining reliable solutions for many cases. Nevertheless, this is not fool-proof and there is no guarantee that a global energy minimum will always be found. Second, and perhaps more importantly, it is crucial to emphasize that the proposed method is a single-determinant method which does not provide a complete description of multi-reference quantum states. As a result, it cannot not accurately describe physical states such as open-shell singlets, which are important for the description of certain chemical ground-states. While this could, in principle, lead to a large deviation between $\Delta^{(\mathrm{HF})}$ and $\Delta^{(\mathrm{FCI})}$, it is not clear how large the deviation will be in practice. Below, we provide empirical evidence for the efficacy of the orbital optimization approach for various small-scale systems, comparing with FCI and DMRG-based numerical approaches.
@@
[…]

### v2

@@ A.2 Application: electronic structure Hamiltonian
Appendix B Numerical limitations of orbital optimization procedure
@@
While the orbital optimization method provides a tractable approach to compute symmetry-aware spectral bounds for large-scale systems, it does have important limitations and challenges which we discuss below. First, as alluded in the main manuscript, we found that the numerical procedure was prone to certain convergence issues arising from the highly non-convex landscape encountered in the orbital optimization procedure. In particular, we found that random initial guesses highly affected final converged energy values. To address this issue, we proposed the pre-orthogonalized one-electron and two-electron integrals in Eqs. (28) and (29). We also found that initializing the $N_{\mathrm{orb}}(N_{\mathrm{orb}}-1)/2$ parameters to zero helped in obtaining reliable solutions for many cases. Nevertheless, this is not fool-proof and there is no guarantee that a global energy minimum will always be found. Second, and perhaps more importantly, it is crucial to emphasize that the proposed method is a single-determinant method which does not provide a complete description of multi-reference quantum states. As a result, it cannot accurately describe physical states such as open-shell singlets, which are important for the description of certain chemical ground-states. While this could, in principle, lead to a large deviation between $\Delta^{(\mathrm{HF})}$ and $\Delta^{(\mathrm{FCI})}$, it is not clear how large the deviation will be in practice. Below, we provide empirical evidence for the efficacy of the orbital optimization approach for various small-scale systems, comparing with FCI and DMRG-based numerical approaches.
@@
N2 [cc-pVDZ] $\Delta_{\mu}/2$ 55.6 55.9
H2O [cc-pVDZ] $\Delta_{\mu}/2$ 42.8 42.9
H2O [cc-pVTZ] $\Delta_{\mu}/2$ 63.44 63.95